# General Place Recognition Survey: Towards Real-World Autonomy

### v1

@@ III-A3 Range Sensor-Related Approaches
Handcraft LPR Methods: ScanContext [39] and ScanContext++ [40] transform LiDAR point clouds into bird-eye-view (BEV) images that encode height in each pixel. Building on this, Wang et al. [41] developed the LiDAR IRIS, enhancing rotation invariance using the LoG-Gabor filter. Ring++ [42] utilizes Radon and Fourier transforms on BEV images for improved representation. In BTC [43], key points of a point cloud are projected onto planes to form triangles whose side lengths serve as triangle descriptors.
Data-Driven LPR Methods: The shift towards feature learning in VPR is similarly observed in LPR, where deep learning methods are increasingly favored. Employing neural networks, including CNNs and Transformers, on point cloud data demands the formulation of specialized network architectures for point clouds or the implementation of extra preprocessing steps to align the data with the original network designs.
The initial advancements in LPR are point-based methods, exemplified by PointNet [37] and PointSift [44], which directly process point clouds to extract features without quantization. PointNetVLAD [45] pioneered this direction by combining PointNet [37] with NetVLAD [25] as the description computation. LPD-Net [46] utilized a graph-based approach for spatial distribution analysis; and SOE-Net [47] proposed an orientation embedding based on PointSift [44] and features are enhaned with self-attention; However, the PointNet structure struggles with orientation invariance, showing reduced performance on significantly rotated point clouds [46]. Addressing this, RPR-Net [48] leveraged SPRIN [49] to derive rotation-invariant features, demonstrating promising outcomes. Nonetheless, a critical drawback of point-wise methods is their escalating computational complexity with increasing point count, posing challenges for real-time application efficiency.
@@
Rather than directly manipulating points within neural networks, two alternative categories of LPR methods utilizes voxelization [50, 51, 52] and projection-based techniques [53, 54, 55]. These methods transform point clouds into 3D voxels and 2D grids respectively, serving as a preparatory phase prior to network input. Regarding the former category, MinkLoc3D [50] employs sparse 3D convolution on voxelized point clouds for place descriptor extraction, achieving substantial enhancements over PointNet-based techniques. Its successor, MinkLoc3D-SI [51], incorporates spherical coordinates and intensity data for each 3D point. Besides leveraging the sparse convolution, LoGG3D-Net [52] additionally introduces a local consistency loss that steers the network to consistently learn local features during revisits.
Projection-based methods come in multiple variations. Cylindrical projection transforms point cloud rotations into translations on a 2D image, and thus is yaw-invariant for subsequent convolutional operations. OverlapNet [21] leverages this method for point cloud pre-processing, generating multi-channel images that encapsulate range, intensity, normals, and semantic information. OverlapTransformer [56] extends this approach, applying transformer networks for image analysis. RINet [57] advanced the model by leveraging semantics and geometric features and attention mechanism to build a discriminative global descriptor. Spherical projection, utilized in the SphereVLAD series [53, 58, 54], offers $3$-DoF rotation invariance, crucial for reliable 3D coordinate encoding. Its successor, SphereVLAD++ [58], applied attention mechanism on features after the spherical projection. There are projection methods: DiSCO [59] proposed a differentiable ScanContext-like representation using the polar projection, while BEVPlace [55] transforms point clouds into BEV images and designed the corresponding rotation-invariant network.
@@
Data-Driven RPR Methods: RPR techniques predominantly utilize polar and Cartesian images derived from Radar measurements. Kidnapped Radar [60] leverages a CNN backbone to process polar images for feature extraction. AutoPlace [61] enhances accuracy by employing Doppler measurements to eliminate moving objects and applies a specialized network to encode Radar point clouds, integrating spatial and temporal dimensions, and further refines matches using Radar Cross Section histograms. mmPlace [62] designed a rotating single-chip Radar platform to enhance the perception FoV. Additionally, advancements in RPR have been made through exploring sequence matching [63], self-supervised fusion [64], and data augmentation [65] strategies to enhnace PR precision.
@@ III-B High-Level Representations in Various Formats
High-level representations offer a sophisticated abstraction of a place’s layout, distinguishing themselves from low-level, sensor-dependent models by focusing on advanced constructs like graph structures and embedding vectors. Graphs are adopted to modal implicit relationship among objects, while embeddings can be understood as a high-level feature vectors that are directly processed by deep learning models. An interesting property is that a set of same-size embeddings can be concatenated, regardless that they are generated from images, point clouds or text. This section aims to detail the construction of high-level representations and examine their application in existing PR works, unveiling their evolution and role in boosting PR performance.
III-B1 Graph
Graphs, including object-based graphs and 3D scene graphs, have recently emerged as a powerful representation of a place. A graph is a mathematical structure that is used to model pairwise relations between objects [67]. It consists of vertices (also called nodes) and edges, where the vertices represent the objects and the edges represent the connections or relationships between them. In the context of PR, these entities can be features, landmarks, or regions, and the edges can denote spatial or topological relations. Graph representations offer several advantages, including robustness to viewpoint changes, occlusions, and dynamic scenes. They can also store semantic labels [66] to enhance PR accuracy.
Recent studies [68, 69, 70, 66, 71] have introduced various graph models to depict places and environments. Co-visibility graphs depict the relationships between landmarks and the different viewpoints from which sensors observe these landmarks [68]. Kong et al. [69] constructed the semantic graph which abstracts object instances and their relative position. LOCUS [70] employed the spatio-temporal higher-order pooling graphs to merge various features including appearance, topology, and temporal links for a unified scene depiction. The topological semantic graph is designed to enable the goal-directed exploration [71]. The Hydra [66] system construct 3D scene graph to represent places with a hierarchical graph structure from low-level metric maps to high-level object semantics. Comparing the similarity of two graphs become the key challenge in PR. Related solutions including graph kernel formulation [68], inner product of features from graph neural network [69], Euclidean distance after feature pooling [70], and hierarchical descriptor matching [66] have been proposed.
@@ III-B2 Embeddings
Implicit embeddings, often referred to as latent codes in the literature, differ from the global descriptors discussed in Section III-A. These embeddings are outputs from intermediate layers of neural networks that are not specifically designed for PR. Although they have been seldom mentioned in previous literature, we have found that implicit embeddings demonstrate significant potential for PR. The RNR-Map [72], for instance, leverages image-based embeddings to construct a 2D grid map that is not only abstract and descriptive but also generalizable and capable of real-time performance. These embeddings, originating from image observations, enable image rendering via volume rendering techniques, yielding a visually rich map useful for localization and navigation. Given their capacity to retain the appearance of places, these embeddings can be directly applied in PR through cross-correlation.
Recent works have explored image-language descriptors for PR, leveraging Visual Language Models (VLM) to link visual and textual information. The CLIP method [73], for instance, transforms images and natural language prompts into a shared embedded space, facilitating comparison between visual and textual representations [74, 75]. LEXTS [74] further integrates CLIP features with topological graph nodes for indoor, room-level PR, employing cosine similarity to gauge distances between image and room text encodings. These approaches enable the language-based data fusion in PR, demonstrating with enhanced resilience to changes in conditions, viewpoints, and overall generalizability. Moreover, the flexibility of embeddings surpasses traditional string comparison [76], mitigating the noise often present in LLM outputs. For instance, LLM may label a single location with synonyms like “corridor” and “hallway”, despite being different terms, share similar embedded representations.
@@ III-C Summary
The methodologies employed to accurately represent places have undergone considerable evolution. Initially, these methods relied heavily on handcrafted techniques necessitating extensive engineering and domain-specific knowledge. The field then shifted towards leveraging pre-trained neural network models for the task of feature extraction, paving the way for the development of innovative end-to-end solutions such as NetVLAD, specifically designed for PR tasks. High-level representations provide a new aspect to place abstraction, demonstrating robustness against conditional and viewpoint changes, along with enhanced flexibility in multi-modal information fusion. Recent advancements in ViTs and foundation models, which are pre-trained on extensive datasets, have significantly enhanced zero-shot generalization ability in feature extraction across various domains. The next section presents methods for determining whether a place has been previously visited, utilizing proper representation and considering various practical factors relevant to robotic navigation.
@@
Place Modeling: semantics [33], domain transfer [77], event cameras [36]
Place Matching with Sequences: sequence matching [78], dynamic time warping [79]
@@
Loss Functions: rotation triplet [53], angular [85], divergence [86], soft binary-cross entrpy [57]
@@
IV Recognizing the Right Place Aginst Challenges
As we stated in Section. II, the primary challenges for place recognition can be categorized into five types: (1) Appearance Change, (2) Viewpoint Difference, (3) Generalization Ability, (4) Efficiency, and (5) Uncertainty Estimation. We will investigate them and review existing solutions separately.
Figure 7: Two typical solutions to appearance change. (1) The i3dLoc system [54] transforms panoramic images from the RGB color domain to the depth image domain, facilitating data association with LiDAR measurements. This method falls under the category of place modeling, as it explicitly models the 3D appearance of environments. (2) SeqNet [104] is categorized as a sequence-based method, as it involves comparing sequences of query images with sequences from a database of images.
@@ IV-A Appearance Change
Appearance changes can introduce recognition failures for the same areas and perceptual aliasing where different places generate similar observations. Two types of appearance changes are commonly presented in PR:
  • Conditional changes, contains the appearance changes caused by environmental conditions, such as illumination, weather, and seasons. This change will mainly affect visual observations over time, causing perceptual aliasing and wrong data association.
  • Structural changes, contains the dynamic objects, geometric transformations, and landform changes over short-term or long-term navigation. These changes can be due to natural phenomena, such as seasonal variations and weather conditions, or human activities, including construction and urban development.
Two categories of solutions with complementary strengths exist to address appearance changes [105]: (1) place modeling, which aims to extracing condition-invariant features to represent a place, and (2) place mathcing with sequences, which estimates the place similarity with a sequence of observations.
@@ IV-A1 Place Modeling
IV-A2 Place Mathcing with Sequences
Since in robot navigation, data is always captured in sequence, Milford et al. established a benchmark [78] based on image sequences instead of single frame such as FAB-MAP. It aggregates scores across defined paths to ascertain the most accurate match, enhancing PR performance significantly, even with basic image intensity normalization. SeqSLAM’s methodology has inspired a vast array of research [108, 109, 110, 111, 104], fostering advancements in addressing more intricate challenges. The original SeqSLAM design presents several limitations. A significant hurdle is its computational demand, which intensifies with the expansion of the reference database and the image sequence’s length. FastSeqSLAM [108] addresses this by employing an approximate nearest neighbor (ANN) search, offering a more efficient alternative to exhaustive search methods. Moreover, Bampis et al. [109] enhance sequence matching with a BoW that includes a temporal consistency filter for improved accuracy. A further complication arises from SeqSLAM’s assumption of constant sensor velocity, making it vulnerable to velocity fluctuations. The efficacy of sequence-based recognition also heavily relies on the quality of single-image descriptors, which may falter under drastic visual changes.
@@ IV-B Viewpoint Difference
PR solutions to address viewpoint differences can be categorized into three primary groups: geometric, appearance-based, and hybrid approaches, each leveraging different input types. Geometric methods, predominantly utilized in LPR systems, include innovations like OverlapNetTransformer [56], which transforms yaw differences in point clouds into translational differences on images via cylinder projection, addition to translation-invariant CNN to extract features. RPR-Net[48] achieves rotation-invariant LPR by utilizing SPRIN[112] rotation-invariant local features and geometry constrains which are consistent within different viewpoints. Appearance-based methods aim to identify visual cues immune to viewpoint shifts. Semantics-aware PR methods, as proposed by Garg et al., facilitate PR across inverse directions. Techniques such as MixVPR [82] utilize global descriptors with attention-weighted patch tokens and isotropic MLP stacks, respectively, to maintain consistent performance despite viewpoint changes. Patch-NetVLAD [83] focuses on extracting patch-level features for global descriptor computation, enhancing viewpoint invariance through a multi-scale patch feature fusion strategy. Hybrid methods, like AutoMerge [84], incorporating both point-based (geometry) and projection-based (appearance) feature extraction. This combination addresses the issue caused by translation and orientation disparities, offering a robust framework for PR under varied viewpoints.
@@ IV-C Generalization Ability
Generalization ability refers to a system’s capacity to recognize places that it is not included during training, which is especially challenging if the unseen environment has drastic variations from the training data [113]. However, infinite real-world environments make it impossible to train data covering all types of scenarios. Enhancing the generalization ability of robots is essential for their autonomous operation, particularly in applications such as autonomous navigation, where they must adeptly navigate through entirely novel environments.
In Section III, the transformative impact of deep learning on place representation is thoroughly explored, highlighting the integration of pre-trained CNN architectures [114], adaptable and trainable frameworks [25], attention mechanisms [58], and ViTs [31]. A notable trend is the enhancement of feature extraction generalization across diverse domain datasets through the augmentation of network capacity (e.g., increased model depth and a greater number of trainable parameters) and the expansion of training data scale. However, as explained in Section IV-D, this augmentation in network size introduces significant efficiency challenges to real-time applications.
Semantics that encode high-level human knowledge can help PR for better inference. PSE-Match [86] designed a semantic feature encoding to extracts the different types of semantics (tree, building, etc.). Paolicelli et al. [91] combined the visual appearance and semantic context through a multi-scale attention module for robust feature embedding. Visual-LiDAR fusion, as demonstrated in AdaFusion [90] and MinkLoc++ [35], enhances generalization capability in PR, surpassing what a single sensor alone can achieve.
@@
Loss functions play a crucial role in training generalized PR models. Triplet loss, a well-established metric, aims to reduce the distance between query-positive pairs while increasing the separation from negative pairs [25, 45]. To enhance orientation invariance, Yin et al. [53] proposed the rotation triplet loss. Angular loss, leveraging cosine similarity, offers robustness against similarity transformations and is effective in handling spatial discrepancies [85]. However, these metrics predominantly concentrate on the relational distances within and between clusters, somewhat overlooking the multifaceted nature of PR challenges. Alternatively, divergence loss, introduced by Yin et al. [86], targets varying semantic structures. Meanwhile, classification-based strategies like RINet [57] formulate PR as classification problem, presenting a soft binary-cross entropy loss for the model training.
Rather than expanding models’ capacity to address PR by training on extensive datasets like Anyloc, lifelong or continual learning chooses to incrementally update agents’ knowledge during missions while utilizing a standard-sized model. AirLoop [87] proposed two loss fucntions to protect the model from catastrophic forgetting when being adapted to a new domain: (1) the relational memory-aware synapses loss, which assigns an importance weight to each model parameter, thus regularizing the parameters’ adjustments throughout the training process; (2) the relational knowledge distillation (RKD) loss, designed to preserve the embedding space structure. In contrast to the RKD loss, InCloud [88] designed a higher-order angular distillation loss. Fig. 10 visulized the key insight of InCloud. HM4 [89] developed a Hidden Markov Model featuring a dual-tier memory management system. Additionally, Bio-SLAM [17] took inspiration from human memory mechanisms to propose a lifelong PR system capable of incremental learning across new domains.

### v2

@@ III-A3 Range Sensor-Related Approaches
Handcraft LPR Methods: Early approaches such as ScanContext [49] and ScanContext++ [50] encode LiDAR point clouds into bird’s-eye-view (BEV) images, where pixel intensities represent height information. Building on this foundation, Wang et al. [51] enhanced rotation invariance in LiDAR PR through LiDAR IRIS, leveraging LoG-Gabor filters for improved feature extraction. A work extended ScanContext for map matching with OpenStreetMap data [52], while Ring++ [53] applied Radon and Fourier transforms to BEV images, enhancing feature representation through frequency domain analysis. Recent innovations such as BTC [54] adopt geometric descriptors by projecting key points onto planes and deriving triangular features from their spatial relationships.
Data-Driven LPR Methods: The transition to data-driven feature learning observed in VPR has extended to LRP, motivated by neural networks’ capacity to learn complex geometric relationships from raw sensor data. However, implementing conventional architectures like CNNs and Transformers for unstructured 3D point clouds necessitates either specialized network designs (e.g., point-voxel transformers) or geometric preprocessing to reconcile irregular point distributions with structured computational paradigms.
Early advancements in LPR emerged through point-based methods, exemplified by PointNet [47] and PointSift [55], which process raw point clouds without voxelization. PointNetVLAD [56] established a foundational framework by merging PointNet’s geometric feature extraction with NetVLAD’s descriptor learning. Subsequent efforts expanded this paradigm: LPD-Net [57] introduced graph-based neighborhood modeling to encode spatial relationships, while SOE-Net [58] augmented local features through self-attention mechanisms guided by PointSift-derived orientation embeddings [55]. A critical limitation of PointNet-based approaches, however, lies in their sensitivity to viewpoint rotations, which degrades performance under large orientation changes [57]. To address this, RPR-Net [59] integrated the rotation-invariant SPRIN backbone [60], demonstrating improved robustness. Despite progress, point-wise methods face scalability constraints due to quadratic complexity growth with point cloud density, limiting their practicality for real-time deployment.
@@
Rather than directly manipulating points within neural networks, two alternative categories of LPR methods utilizes voxelization [61, 62, 63] and projection-based techniques [64, 65, 66]. These methods transform point clouds into 3D voxels and 2D grids respectively, serving as a preparatory phase prior to network input. Regarding the former category, MinkLoc3D [61] employs sparse 3D convolution for feature extraction. Its successor, MinkLoc3D-SI [62], incorporates spherical coordinates and intensity data for each 3D point. Besides leveraging the sparse convolution, LoGG3D-Net [63] additionally introduces a local consistency loss that steers the network to consistently learn local features during revisits.
Initially, both point-based and voxel-based methods struggle to handle large viewpoint differences. However, by effectively utilizing point cloud projections, these differences can be more easily mitigated through advanced image processing techniques. Projection-based methods in LPR vary in approach. Cylindrical projection, which converts point cloud rotations to translations in a 2D image, provides yaw-invariance for convolutional processing. OverlapNet series [20, 67] utilizes this method, creating multi-channel images with range, intensity, normals, and semantic information, redefining PR as a classification task based on scan overlap. RINet [68] further advances it using semantic and geometric features with attention mechanisms. Spherical projection, as used in SphereVLAD series [64, 69], offers 3-DoF rotation invariance, essential for consistent 3D coordinate encoding. Other projection methods include DiSCO [70], which applies a differentiable ScanContext-like representation using polar projection, and BEVPlace [66], which transforms point clouds into BEV images with a rotation-invariant network design.
Re-ranking mechanisms have also been incorporated into LPR. Unlike Patch-NetVLAD, SpectralGV [71] expands this approach by using Spectral Matching to calculate matching confidence. TReR [72] introduces a transformer-based re-ranking method that relies solely on global descriptors, bypassing the need for local features.
@@
Data-Driven RPR Methods: RPR techniques predominantly utilize polar and Cartesian images derived from Radar measurements. Kidnapped Radar [73] leverages a CNN backbone to process polar images for feature extraction. AutoPlace [74] enhances accuracy by employing Doppler measurements to eliminate moving objects and applies a specialized network to encode Radar point clouds, integrating spatial and temporal dimensions, and further refines matches using Radar Cross Section histograms. mmPlace [75] designed a rotating single-chip Radar platform to enlarge the FoV. Additionally, advancements in RPR have been made through exploring sequence matching [76], cross-modal data matching with LiDARs [77] and overhead images [78], self-supervised fusion [78], and data augmentation [79] strategies to enhance RPR.
@@ III-B High-Level Representations in Various Formats
High-level representations in PR provide a semantic abstraction of environmental structure, diverging from low-level feature-based approaches that rely on raw geometric primitives or sensor-specific measurements. They capture topological relationships and contextual attributes through structured formats such as graphs or learned embeddings. Graph-based methods model spatial-semantic interactions (e.g., object adjacency), while neural embeddings project multi-modal inputs (images, point clouds, text) into unified descriptor spaces, enabling cross-domain compatibility through fixed-dimensional vector concatenation. This abstraction not only decouples PR systems from sensor dependencies but also facilitates hierarchical reasoning about place identity. In this section, we analyze the design principles of high-level representations, survey their implementation in modern PR frameworks, and unveil their evolution and role in boosting PR performance.
III-B1 Graphs
Graphs, including object-based graphs and 3D scene graphs, have recently emerged as a powerful representation of a place. A graph is a mathematical structure that is used to model pairwise relations between objects. It consists of vertices (also called nodes) and edges, where the vertices represent the objects and the edges represent the connections or relationships between them. In the context of PR, these entities can be features, landmarks, or regions, and the edges can denote spatial or topological relations. Graph representations offer several advantages, including robustness to viewpoint changes, occlusions, and dynamic scenes. They can also store semantic labels [80] to enhance PR accuracy.
Recent studies [81, 82, 80, 83, 84, 85] have introduced various graph models to depict places and environments. Co-visibility graphs depict the relationships between landmarks and the different viewpoints from which sensors observe these landmarks [86]. Kong et al. [81] constructed the semantic graph which abstracts object instances and their relative position. LOCUS [82] employed the spatio-temporal higher-order pooling graphs to merge features including appearance, topology, and temporal links for a unified scene depiction. The topological semantic graph is designed to enable the goal-directed exploration [83]. The Hydra [80] system construct 3D scene graph to represent places with a hierarchical graph structure from low-level metric maps to high-level object semantics. Comparing the similarity of two graphs become the key challenge in PR. Related solutions including graph kernel formulation [86], inner product of features from graph neural network [81], Euclidean distance after feature pooling [82], and hierarchical descriptor matching [80] have been proposed.
@@ III-B2 Embeddings
Implicit embeddings differ fundamentally from global descriptors (Section III-A): while descriptors are explicitly trained for PR, embeddings are generic latent codes from networks optimized for unrelated tasks. Despite lacking PR-specific training, these embeddings encode environmental semantics through their native objectives. Although the 2D grid embeddings of RNR-Map [87] were trained from image observations via volume rendering techniques, they inherently capture spatial-appearance features usable for PR by cross-correlation. Their dual utility stems from architectural constraints: the same latent codes enabling rendering also structure information into a format amenable to place matching, proving that task-agnostic embeddings can achieve PR.
Recent works have explored image-language descriptors for PR, leveraging visual-language models (VLM) to link visual and textual information [88, 89, 90]. The CLIP model [91] aligns images and text through contrastive learning, projecting both modalities into a shared embedding space where semantically related visual and textual representations exhibit high similarity. LEXIS [88] further integrates CLIP features with topological graph nodes for indoor, room-level PR, employing cosine similarity to gauge distances between image and room text encodings. These approaches enable the language-based data fusion in PR, demonstrating with enhanced resilience to changes in conditions, viewpoints, and overall generalizability. Embeddings surpass traditional string comparison [92] by resolving synonym discrepancies despite differing strings. For example, synonyms like “corridor” and “hallway” yield similar embeddings through semantic alignment, avoiding the brittleness of exact text matches
III-B3 Others
Several novel representations also present potential for PR. For instance, researchers have used view synthesis methods [93, 94, 95] to increase the density of database images. Qi et al. [96] utilized GS to store multi-modal data (including images and point clouds), eliminating the need for separate feature extraction modules for different sensors. Moreover, Brachmann et al. [97] developed a regression network that implicitly learns place representation by encoding the whole scene as parameters of a neural network, offering advantages in storage and privacy preservation. This method is categorized as scene coordinate regression [98, 99]: regressing 3D-2D correspondences from the query image to the global scene.
@@ III-C Summary
Place representation methodologies have evolved significantly. Early approaches relied on handcrafted features requiring substantial domain expertise. The field transitioned to pre-trained neural networks for feature extraction, enabling end-to-end solutions like NetVLAD tailored for PR. High-level representations introduced robustness to viewpoint and environmental variations while enabling multimodal fusion. Recent vision transformers (ViTs) and foundation models, trained on large-scale datasets, further advanced zero-shot generalization across domains. The following section discusses place matching strategies that leverage these representations, addressing practical challenges in robotic navigation.
@@
Place Modeling: semantics [44], domain transfer [100], event cameras [46, 101]
Place Matching with Sequences: sequence matching [102], dynamic time warping [103], sequential descriptor [104]
@@
Loss Functions: rotation triplet [64], angular [108], divergence [109], soft binary-cross entropy [68], large margin cosine [30]
@@
IV Recognizing the Right Place Against Challenges
As we stated in Section. II, the primary challenges for PR can be categorized into five types: (1) Appearance Change, (2) Viewpoint Difference, (3) Generalization Ability, (4) Efficiency, and (5) Uncertainty Estimation. We will investigate them and review existing solutions separately.
Figure 7: Two typical solutions to appearance change. (1) The i3dLoc [65] transforms panoramic images from the RGB color domain to the depth image domain, facilitating data association with LiDAR measurements. This method falls under the category of place modeling, as it explicitly models the 3D appearance of environments. (2) SeqNet [104] is a sequence-based method, as it employs learned sequential descriptors to compare image sequences.
@@ IV-A Appearance Change
Two types of changes are commonly presented in PR:
  • Conditional changes, contains the appearance changes caused by environmental conditions, such as illumination, weather, and seasons. This change will mainly affect cameras’ observations over time, causing perceptual aliasing and wrong data association.
  • Structural changes, contains the dynamic objects, geometric transformations, and landform changes over short-term or long-term navigation. These changes can be due to natural phenomena, such as seasonal variations and weather conditions, or human activities, including construction and urban development. They mainly affect measurements of range sensors.
Two categories of solutions with complementary strengths exist to address appearance changes [128]: (1) place modeling, which aims to extracting condition-invariant features to represent a place, and (2) place mathcing with sequences, which estimates the place similarity with a sequence of observations.
@@ IV-A1 Place Modeling
IV-A2 Place Matching with Sequences
Researchers have explored sequential information for PR, capitalizing on the inherent temporal continuity of robot navigation trajectories. However, this type of approach fundamentally requires that query and database paths share sequential overlap, i.e., temporal continuity in visited locations. Milford et al. introduced SeqSLAM [102], which replaces single-image matching with sequence alignment: it aggregates similarity scores across consecutive frames using basic normalized image descriptors, significantly improving robustness to environmental variations.
While foundational, SeqSLAM faces two key limitations: (1) Computational complexity scaling with database size and sequence length, addressed via approximate nearest neighbor search in FastSeqSLAM [131]; (2) Sensitivity to velocity variations, partially mitigated by Bampis et al.’s temporally filtered BoW approach [132]. To balance recall and efficiency, SeqNet [104] introduced a hierarchical solution using a learned sequential descriptor to generate candidate matches, bypassing exhaustive database searches while suppressing false positives through subsequent sequence verification. This contrasts with methods such as [133, 134] that proposed the coarse-to-fine matching strategies.
@@ IV-B Viewpoint Difference
PR solutions to address viewpoint differences can be categorized into three primary groups: geometric, appearance-based, and hybrid approaches, each leveraging different input types. Geometric methods, predominantly utilized in LPR systems, include innovations like OverlapNetTransformer [67], which transforms yaw differences in point clouds into translational differences on images via cylinder projection, addition to translation-invariant CNN to extract features. RPR-Net [59] achieves rotation-invariant LPR by utilizing SPRIN [135] rotation-invariant local features and geometry constrains which are consistent within different viewpoints. Appearance-based methods aim to identify visual cues immune to viewpoint shifts. Semantics-aware PR methods, as proposed by Garg et al., facilitate PR across inverse directions. Techniques such as MixVPR [28] utilize global descriptors with attention-weighted patch tokens and isotropic MLP stacks, respectively, to maintain consistent performance despite viewpoint changes. EigenPlaces [136] proposed a novel method to train the network on images from different perspectives. Patch-NetVLAD [27] focuses on extracting patch-level features for global descriptor computation, enhancing viewpoint invariance through a multiscale patch feature fusion strategy. Hybrid methods, like AutoMerge [107], incorporating both point-based (geometry) and projection-based (appearance) feature extraction. This combination addresses the issue caused by translation and orientation disparities, offering a robust framework for PR under varied viewpoints.
@@ IV-C Generalization Ability
Generalization ability denotes a PR algorithm’s capacity to identify locations under environmental conditions absent from its training data. This capability proves critical when handling domain shifts, such as variations in visual appearance (e.g., day/night cycles), structural layouts (e.g., cross-city topological differences), or perceptual conditions (e.g., perspective versus panoramic imaging). For instance, an autonomous robot must reliably operate in scenarios ranging from snow-covered rural roads to densely structured urban canyons, despite never encountering these specific configurations during training. This section introduces four categories of solutions to enhance generalization: domain generalization, domain adaptation, loss functions, and lifelong learning.
Domain Generalization trains models to maintain robustness on unseen target domains by minimizing distributional discrepancies between training and testing environments [137]. Building on advancements in place representation such as pre-trained CNNs (e.g., VGG [138]), adaptable architectures like NetVLAD [26], attention mechanisms [69, 139], and Vision Transformers (ViTs) [32], the methods significantly enhance the cross-domain feature transfer. Another solution is the data manipulation, which enhances training diversity in PR by simulating domain shifts through: geometric transformations (e.g., rotation and resize), image erasing to simulate occlusions, and photometric synthesis of lighting/weather effects [140].
Domain Adaptation addresses scenarios where unlabeled target domain data is available during training, enabling model adaptation from a labeled source domain to the target distribution. Knights et al. [141] explored test-time adaptation under domain shift between training and test distributions without ground-truth labels. They proposed GeoAdapt, using geometric consistency to generate pseudo-labels and retrain the model for target domain adaptation. Semantics that encode high-level human knowledge can enhance PR’s generalization [109, 115]. PSE-Match [109] separately extracts features from point cloud in different semantics (tree, building, etc.), enabling the descriptors more stable and consistent across different environments. Visual-LiDAR fusion, as demonstrated in AdaFusion [113] and MinkLoc++ [114], enhances generalization capability in PR, surpassing what a single sensor alone can achieve.
@@
Loss Functions also play a crucial role in training generalized PR models. Triplet loss, a well-established metric, aims to reduce the distance between query-positive pairs while increasing the separation from negative pairs [26, 56]. To enhance orientation invariance, Yin et al. [64] proposed the rotation triplet loss. Angular loss, leveraging cosine similarity, offers robustness against similarity transformations and is effective in handling spatial discrepancies [108]. However, these metrics predominantly concentrate on the relational distances within and between clusters, somewhat overlooking the multifaceted nature of PR challenges. Alternatively, divergence loss was introduced in [109]to target varying semantic structures. Meanwhile, classification-based strategies like RINet [68] formulate PR as classification problem, presenting a soft binary cross entropy loss for the model training. Berton et al. [30] introduced the Large Margin Cosine Loss to bypass computationally expensive negative sample mining, thereby enabling scalable training on large-scale datasets.
The inherent variability of real-world environments renders exhaustive training data coverage fundamentally unattainable. Rather than pursuing unsustainable model scaling through massive datasets [32], a paradigm shift emerges: Lifelong Learning [142] (also termed continual learning) emulates biological learning systems by enabling autonomous agents to incrementally integrate new domain knowledge during operational deployment while maintaining fixed model capacity. AirLoop [110] proposed two loss fucntions to protect the model from catastrophic forgetting when being adapted to a new domain: (1) the relational memory-aware synapses loss, which assigns an importance weight to each model parameter, thus regularizing the parameters’ adjustments throughout the training process; (2) the relational knowledge distillation (RKD) loss, designed to preserve the embedding space structure. In contrast to the RKD loss, InCloud [111] designed a higher-order angular distillation loss. Fig. 10 visulized the key insight of InCloud. CCL [143] identified the limitation of using triplet loss in InCloud and instead applied contrastive loss to encourage the model to extract more generalizable features. There are approaches, considering the real-world robotic applications, discussed in Section V-D.